# Robust Fine-Tuning from Non-Robust Pretrained Models: Mitigating Suboptimal Transfer With Adversarial Scheduling

### v1

@@ 1 Introduction
Fine-tuning pretrained models (backbones) on a downstream task is the standard workflow in machine learning, spanning natural language processing (Koroteev, 2021) and computer vision (Goldblum et al., 2023). This workflow offers clear benefits: (i) reusing a single foundation model across tasks (Devlin et al., 2019), (ii) faster convergence, (iii) better generalization than training from scratch (Yosinski et al., 2014), and (iv) reduced computation (Weiss et al., 2016), especially when labelled data is scarce (Pan & Yang, 2010).
@@
When fine-tuning on a downstream task with a robust objective results in near-random performance, the benefits of using a pretrained model are diminished. This raises the question: do standard pretrained models fail to offer a beneficial initialization for training a robust model? In this study, we explore the challenges associated with robust fine-tuning using standard pretrained models and propose a novel approach to address these obstacles. In contrast to standard fine-tuning, where model adaptation to the downstream task occurs immediately, our study reveals that in robust fine-tuning, task adaptation is delayed until later epochs. This delay seems to eventually lead in suboptimal transfer and we observe that the duration of the delay correlates negatively with transfer performance.
@@
Figure 2: Epsilon-Scheduling
To better evaluate the fine-tuned models, we introduce a complementary evaluation metric, dubbed expected robustness. This proposed metric evaluates the expectation of the model accuracy across the full perturbation range from zero (clean accuracy) to the target robustness threshold. The expected robustness provides a concise yet comprehensive measure of the accuracy–robustness trade-off, grounded in a practical threat model. We demonstrate that it offers valuable insights for model selection. Under this metric, Epsilon-Scheduling consistently outperforms the standard robust-finetuning, even when worst-case robustness at the target threshold is similar or lower.
Summary of Contributions Our main contributions are: (i) We show that robust fine-tuning from non-robust backbones often leads to suboptimal transfer, even at small perturbation strengths, where performance can fall significantly below standard fine-tuning. (ii) We find that robust finetuning results in task adaptation delay compared to standard finetuning and that this delay strongly correlates with suboptimal transfer. (iii)We propose Epsilon-Scheduling, a two-hinge linear schedule for adjusting the training perturbation strength, which effectively mitigates the challenges associated with optimizing adversarial loss. (iv) We introduce expected robustness, a new evaluation metric capturing the full accuracy–robustness trade-off, and report performance using this metric for the first time. (v) Through extensive experiments, we show that Epsilon-Scheduling consistently prevents suboptimal transfer and improves expected robustness across both moderate ($\nicefrac{{4}}{{255}}$) and high ($\nicefrac{{8}}{{255}}$) perturbation regimes.
@@ 2 Related work
Adversarial Robustness in Transfer Learning with Robust-FineTuning There are two main ways to achieve adversarial robustness in Transfer Learning: Robust Distillation (Goldblum et al., 2020; Dong et al., 2024) and Robust Fine-Tuning. Prior works on RFT focus on robust backbones (Liu et al., 2023; Xu et al., 2024; Hua et al., 2024). TWINS (Liu et al., 2023) employs two networks with shared parameters to separately track pretraining and downstream batch statistics. AutoLoRA (Xu et al., 2024) disentangles natural and adversarial objectives using a LoRA branch for the former and a robust pretrained extractor for the latter, though it relies on TRADES loss (Zhang et al., 2019), which is harder to scale than standard adversarial training (Madry et al., 2018). RoLi (Hua et al., 2024) preserves robustness by initializing the classifier head via adversarial linear probing before performing RFT. All these approaches assume robust pretrained features. In contrast, we are the first to propose an RFT method targeting non-robust backbones.
@@
Tuning Perturbation Strength in Adversarial Training Adapting the perturbation strength during training has been explored in various forms. Early work used a linear ramp-up in Interval Bound Propagation (Gowal et al., 2018). Ding et al. (2020) connected margin maximization to minimal adversarial loss, motivating adaptive, sample-specific perturbation strengths, though such instance-wise selection (Balaji et al., 2019) is computationally costly. They also proposed PGDLS (PGD with Linear Scaling), which linearly increases perturbation strength but shows gains only at large perturbations ($\nicefrac{{24}}{{255}}$). Other strategies include sampling the perturbation strength from a Beta distribution (Chamon & Ribeiro, 2020) and curriculum schemes that gradually increase the number of attack steps (Cai et al., 2018). Pang et al. (2021) reports that linear warmup provides limited gains in ResNets, whereas Debenedetti et al. (2023) finds that it improves both clean and robust accuracy in vision transformers. Unlike prior works, which apply to adversarial training from scratch, our study is on transfer learning. Our formulation generalizes linear warmup, and we show that the benefits consistently hold across tasks and architectures, including ResNets, using the new expected robustness metric.
@@
Adversarial Defense Evaluation Standard evaluation (Croce et al., 2021) compares clean and robust accuracy at a target perturbation strength under strong attacks or ensembles (Carlini & Wagner, 2017; Madry et al., 2018; Croce & Hein, 2020b; Cinà et al., 2025), yet it obfuscates what happens at intermediate perturbation strengths. A related recommendation is to check that accuracy decreases with stronger perturbations (Carlini et al., 2019), but this test is unquantified and serves only as an informal validation (Debenedetti et al., 2023). In contrast, our notion of expected robustness formalizes this decrease and interpolates between clean and worst-case performance at a specific target perturbation strength. Another class of metrics interpolates between worst-case and average-case robustness (Rice et al., 2021; Li et al., 2021), the latter defined against random or natural perturbations (Hendrycks & Dietterich, 2019; Han et al., 2024). However, this approach does not capture the trade-off between clean and worst-case performance.
@@ 4 Characterizing Suboptimal Transfers in Robust Fine-Tuning
Suboptimal Transfer We apply RFT-fix with target perturbation strengths $\varepsilon_{g}$ in $\nicefrac{{1}}{{255}},\nicefrac{{9}}{{255}}]$ and compare performance (clean accuracy) to standard fine-tuning ($\varepsilon_{g}=0$). While low clean accuracies for RFT-fix from non-robust backbones have been reported on Caltech at $\varepsilon_{g}=8/255$ (Liu et al., 2023; Hua et al., 2024), our experiments reveal a broader picture. Figure 1 shows distinct trends in clean accuracy degradation as $\varepsilon_{g}$ increases. Even for tiny perturbations ($\varepsilon_{g}=\nicefrac{{1}}{{255}}$), accuracy can drop by up to $14\%$ compared to the standard finetuning with $\epsilon_{g}=0$. At the commonly used threshold $\varepsilon_{g}=\nicefrac{{4}}{{255}}$, the minimum drop is $10\%$. In extreme cases, RFT-fix fails to adapt effectively, with clean accuracy falling below $5\%$, making it practically unusable. We refer to this phenomenon as suboptimal transfer, where RFT-fix yields a clean accuracy substantially lower than standard fine-tuning, at times to the point of ineffectiveness or failure.
Insights on Task Difficulty and Backbone Selection  The severity of suboptimal transfer depends on both the backbone and the task, with the task being the more differentiating factor (see Figure Figure 1). Easier tasks (e.g., Caltech, wide variety of objects) exhibit smaller performance drops than more challenging ones (e.g., Aircraft, highly similar categories). As expected, the choice of SWIN as backbone is better than ViT (Hua et al., 2024). However, the notion of “task difficulty" is also model-dependent: (Dogs << Cub for SWIN but Dogs >> Cub for ViT). This demonstrates that task difficulty is a property that emerges from the interaction between model and dataset (Zilly et al., 2019; Jankowski et al., 2025), rather than by features alone (Cao et al., 2024). Our experiments further suggest that with respect to adversarial robustness, the perturbation strength may also play a role: Dogs is easier than Cub for perturbations below $\varepsilon_{g}=\nicefrac{{4}}{{255}}$, but becomes comparable or more difficult at higher strengths.
@@ 5 Epsilon-Scheduling and Expected Robustness
The expected robustness metric also evaluates performance under a more realistic threat model where inputs may or may not be altered. Clean accuracy corresponds to the idealized case where inputs are never perturbed, while robust accuracy at $\varepsilon_{g}$ assumes that all inputs are perturbed with a budget of $\varepsilon_{g}$. In contrast, expected robustness estimates the accuracy when perturbations of size up to $\varepsilon_{g}$ are applied at random (here, uniformly). While our choice of a uniform distribution serves as a good option from a use-case-agnostic perspective, the distribution of perturbations can be tailored to align with the relevant threat model for a specific application.
@@ 6 Experimental Evaluation
Training and Evaluation We follow a similar setup described in Hua et al. (2024). We train for 50 epochs and generate adversarial examples using APGD (Croce & Hein, 2020b) (instead of PGD) with cross-entropy loss as in prior work (Singh et al., 2023; Heuillet et al., 2025), which removes the need for manual tuning. The number of APGD steps is 7 for training and 10 for evaluation (Hua et al., 2024). Results are reported for the models at the end of training because overfitting of the adversarial accuracy is negligible (Figure 6). The evaluation is conducted in the $\ell_{\infty}$-norm, which is the most widely studied norm in the literature (Croce et al., 2021; Ngnawe et al., 2024), using two standard evaluation thresholds: $\varepsilon_{g}=\nicefrac{{4}}{{255}}$ (moderate perturbation) and $\varepsilon_{g}=\nicefrac{{8}}{{255}}$ (high perturbation). For each model, we report the clean accuracy (clean), APGD accuracy (adv.), and the expected APGD accuracy (E. adv.) over the interval $0,\varepsilon_{g}]$. We provide in Appendix B.2 a few additional results for SWIN, with the more expensive evaluation AutoAttack (Croce & Hein, 2020b), a diverse ensemble of four attacks containing untargeted APGD-CE, targeted APGD-DLR, targeted FAB (Croce & Hein, 2020a), and black-box Square Attack (Andriushchenko et al., 2020).
@@ 6.2 Results Analysis
$\varepsilon_{g}=\nicefrac{{4}}{{255}}$ $\varepsilon_{g}=8/255$
@@ 6.3 Ablation and Sensitivity Analysis
We summarize the effect of the hyperparameters $T_{1}$ and $T_{2}$ of Epsilon-Scheduling in Appendix C.1 as follows. (i) $T_{2}$ have the most significant influence with the control of steepness $\left(\nicefrac{{1}}{{(T_{2}-T_{1})}},T_{1}\neq T_{2}\right)$. When $T_{2}$ is close to $T_{1}$, clean accuracy decreases, whereas robust accuracy increases; this eventually leads to suboptimal transfer. This is in line with the motivation for linear warmup in Cao et al. (2024), although they do not study this effect. (ii) Increasing $T_{1}$ increases clean accuracy, up to a threshold beyond which further increasing $T_{1}$ has no apparent effect.
@@ 7 Conclusion
Limitations and Future Work. Although Epsilon-Scheduling yields significant improvements, robustness can still be limited even when clean accuracy is decent, highlighting the potential for future research to further enhance performance. This work opens doors to exploring other scheduling strategies, either heuristic, theoretically motivated, or learning-based. Extending the analysis to other vision tasks, such as detection or segmentation, applying the framework to parameter-efficient methods like LoRA, and investigating whether similar dynamics occur in other modalities, such as natural language processing, remain open questions. Studying these cases may require special considerations such as task-specific losses or hyperparameters.

### v2

# arXiv:2509.23325v2[cs.LG] 28 Feb 2026

@@ 1 Introduction
Fine-tuning pretrained models (backbones) on a downstream task is the standard workflow in machine learning, spanning natural language processing (Koroteev, 2021) and computer vision (Goldblum et al., 2023). This workflow offers clear benefits: (i) reusing a single foundation model across tasks (Devlin et al., 2019), (ii) faster convergence, (iii) better generalization than training from scratch (Yosinski et al., 2014), and (iv) reduced computation (Weiss et al., 2016), especially when labelled data is scarce (Pan and Yang, 2010).
@@
When fine-tuning on a downstream task with a robust objective results in near-random performance, the benefits of using a pretrained model are diminished. This raises the question: do standard pretrained models fail to offer a beneficial initialization for training a robust model? In this study, we explore the challenges associated with robust fine-tuning using standard pretrained models and propose a novel approach to address these obstacles. In contrast to standard fine-tuning, where model adaptation to the downstream task occurs immediately, our study reveals that in robust fine-tuning, task adaptation is delayed until later epochs. This delay seems to eventually lead to suboptimal transfer, and we observe that the duration of the delay correlates negatively with transfer performance.
Figure 2: Epsilon-Scheduling
@@
To better evaluate the fine-tuned models, we introduce a complementary evaluation metric, dubbed expected robustness. This proposed metric evaluates the expectation of the model accuracy across the full perturbation range from zero (clean accuracy) to the target robustness threshold. The expected robustness provides a concise, yet comprehensive measure of the accuracy–robustness trade-off, grounded in a practical threat model. We demonstrate that it offers valuable insights for model selection. Under this metric, Epsilon-Scheduling [^1] Code available at: https://github.com/ngnawejonas/EpsilonScheduling consistently outperforms the standard robust-finetuning, even when worst-case robustness at the target threshold is similar or lower.
Summary of Contributions Our main contributions are: (i) We show that robust fine-tuning from non-robust backbones often leads to suboptimal transfer, even at small perturbation strengths, where performance can fall significantly below standard fine-tuning. (ii) We find that robust finetuning results in task adaptation delay compared to standard finetuning and that this delay strongly correlates with suboptimal transfer. (iii)We propose Epsilon-Scheduling, a two-hinge linear schedule to adjust the training perturbation strength, which effectively mitigates the challenges associated with optimizing adversarial loss. (iv) We introduce expected robustness, a new evaluation metric capturing the full accuracy–robustness trade-off, and report performance using this metric for the first time. (v) Through extensive experiments, we show that Epsilon-Scheduling consistently prevents suboptimal transfer and improves expected robustness across both moderate ($\nicefrac{{4}}{{255}}$) and high ($\nicefrac{{8}}{{255}}$) perturbation regimes.
@@ 2 Related work
Adversarial Robustness in Transfer Learning with Robust-FineTuning There are two main ways to achieve adversarial robustness in Transfer Learning: Robust Distillation (Goldblum et al., 2020; Dong et al., 2024) and Robust Fine-Tuning. Prior works on RFT focus on robust backbones (Liu et al., 2023; Xu et al., 2024; Hua et al., 2024). TWINS (Liu et al., 2023) employs two networks with shared parameters to separately track pretraining and downstream batch statistics. However, Liu et al. (2023) do not apply it to non-robust backbones claiming that ”robust pre-training is indispensable to downstream robustness”. AutoLoRA (Xu et al., 2024) disentangles natural and adversarial objectives using a LoRA branch for the former and a robust pretrained extractor for the latter, though it relies on TRADES loss (Zhang et al., 2019), which is harder to scale than standard adversarial training (Madry et al., 2018). Xu et al. (2024) show that robust pretrainning is necessary for AutoLoRA and register the worst performance for the non-robust pretraining. RoLi (Hua et al., 2024) preserves robustness by initializing the classifier head via adversarial linear probing before performing RFT. (Hua et al., 2024) explicitly argue that robust pretraining is a prerequisite for RoLi by showing that linear probing with a robust objective on a non-robust backbone fails dramatically and therefore does not provide a good initialization. In summary, all these approaches assume robust pretrained features. In contrast, we are the first to propose an RFT method targeting non-robust backbones.
@@
Tuning Perturbation Strength in Adversarial Training Adapting the perturbation strength during training has been explored in various forms. Early work used a linear ramp-up in Interval Bound Propagation (Gowal et al., 2018). Ding et al. (2020) connected margin maximization to minimal adversarial loss, motivating adaptive, sample-specific perturbation strengths, though such instance-wise selection (Balaji et al., 2019) is computationally costly. They also proposed PGDLS (PGD with Linear Scaling), which linearly increases perturbation strength but shows gains only at large perturbations ($\nicefrac{{24}}{{255}}$). Other strategies include sampling the perturbation strength from a Beta distribution (Chamon and Ribeiro, 2020) and curriculum schemes that gradually increase the number of attack steps (Cai et al., 2018). Pang et al. (2021) reports that linear warmup provides limited gains in ResNets, whereas Debenedetti et al. (2023) finds that it improves both clean and robust accuracy in vision transformers. Unlike prior works, which apply to adversarial training from scratch, our study is on transfer learning. Our formulation generalizes linear warmup, and we show that the benefits consistently hold across tasks and architectures, including ResNets, using the new expected robustness metric.
@@
Adversarial Defense Evaluation Standard evaluation (Croce et al., 2021) compares clean and robust accuracy at a target perturbation strength under strong attacks or ensembles (Carlini and Wagner, 2017; Madry et al., 2018; Croce and Hein, 2020b; Cinà et al., 2025), yet it obfuscates what happens at intermediate perturbation strengths. A related recommendation is to check that accuracy decreases with stronger perturbations (Carlini et al., 2019), but this test is unquantified and serves only as an informal validation (Debenedetti et al., 2023). In contrast, our notion of expected robustness formalizes this decrease and interpolates between clean and worst-case performance at a specific target perturbation strength. Another class of metrics interpolates between worst-case and average-case robustness (Rice et al., 2021; Li et al., 2021), the latter defined against random or natural perturbations (Hendrycks and Dietterich, 2019; Han et al., 2024). However, this approach does not capture the trade-off between clean and worst-case performance.
@@ 4 Characterizing Suboptimal Transfers in Robust Fine-Tuning
Suboptimal Transfer We apply RFT-fix with target perturbation strengths $\varepsilon_{g}$ in $\nicefrac{{1}}{{255}},\nicefrac{{9}}{{255}}]$ and compare performance (clean accuracy) to standard fine-tuning ($\varepsilon_{g}=0$). While low clean accuracies for RFT-fix from non-robust backbones have been reported on Caltech at $\varepsilon_{g}=\nicefrac{{8}}{{255}}$ (Liu et al., 2023; Hua et al., 2024), our experiments reveal a broader picture. Figure 1 shows distinct trends in clean accuracy degradation as $\varepsilon_{g}$ increases. Even for tiny perturbations ($\varepsilon_{g}=\nicefrac{{1}}{{255}}$), accuracy can drop by up to $14\%$ compared to the standard finetuning with $\epsilon_{g}=0$. At the commonly used threshold $\varepsilon_{g}=\nicefrac{{4}}{{255}}$, the minimum drop is $10\%$. In extreme cases, RFT-fix fails to adapt effectively, with clean accuracy falling below $5\%$, making it practically unusable. We refer to this phenomenon as suboptimal transfer, where RFT-fix yields a clean accuracy substantially lower than standard fine-tuning, at times to the point of ineffectiveness or failure.
Insights on Task Difficulty and Backbone Selection  The severity of suboptimal transfer depends on both the backbone and the task, with the task being the more differentiating factor (see Figure Figure 1). Easier tasks (e.g., Caltech, wide variety of objects) exhibit smaller performance drops than more challenging ones (e.g., Aircraft, highly similar categories). As expected, the choice of SWIN as backbone is better than ViT (Hua et al., 2024). However, the notion of “task difficulty” is also model-dependent: (Dogs << Cub for SWIN but Dogs >> Cub for ViT). This demonstrates that task difficulty is a property that emerges from the interaction between model and dataset (Zilly et al., 2019; Jankowski et al., 2025), rather than by data features alone (Cao et al., 2024). Our experiments further suggest that with respect to adversarial robustness, the perturbation strength may also play a role: Dogs is easier than Cub for perturbations below $\varepsilon_{g}=\nicefrac{{4}}{{255}}$, but becomes comparable or more difficult at higher strengths.
@@ 5 Epsilon-Scheduling and Expected Robustness
The expected robustness metric also evaluates performance under a more realistic threat model where inputs may or may not be altered. Clean accuracy corresponds to the idealized case where inputs are never perturbed, while robust accuracy at $\varepsilon_{g}$ assumes that all inputs are perturbed with a budget of $\varepsilon_{g}$. In contrast, expected robustness estimates the accuracy when perturbations of size up to $\varepsilon_{g}$ are applied at random (here, uniformly). While our choice of a uniform distribution serves as a good option from a use-case-agnostic perspective, the distribution of perturbations can be tailored to align with the relevant threat model for a specific application. With the uniform distribution, each adversarial strength is weighted equally, however, for example, with a Dirac distribution centered at 0 ($\epsilon_{g}$), it falls back to clean accuracy (worst case robust accuracy).
@@ 6 Experimental Evaluation
Training and Evaluation We follow a similar setup described in Hua et al. (2024). We train for 50 epochs and generate adversarial examples using APGD (Croce and Hein, 2020b) (instead of PGD) with cross-entropy loss as in prior work (Singh et al., 2023; Heuillet et al., 2025), which removes the need for manual tuning. The number of APGD steps is 7 for training and 10 for evaluation (Hua et al., 2024). Results are reported for the models at the end of training because overfitting of the adversarial accuracy is negligible (Figure 6). The evaluation is conducted in the $\ell_{\infty}$-norm, which is the most widely studied norm in the literature (Croce et al., 2021; Ngnawe et al., 2024), using two standard evaluation thresholds: $\varepsilon_{g}=\nicefrac{{4}}{{255}}$ (moderate perturbation) and $\varepsilon_{g}=\nicefrac{{8}}{{255}}$ (high perturbation). For each model, we report the clean accuracy (clean), APGD accuracy (adv.), and the expected APGD accuracy (E. adv.) over the interval $0,\varepsilon_{g}]$. We provide in Appendix B.2 a few additional results for SWIN, with the more expensive evaluation AutoAttack (Croce and Hein, 2020b), a diverse ensemble of four attacks containing untargeted APGD-CE, targeted APGD-DLR, targeted FAB (Croce and Hein, 2020a), and black-box Square Attack (Andriushchenko et al., 2020).
@@ 6.2 Results Analysis
$\varepsilon_{g}=\nicefrac{{4}}{{255}}$ $\varepsilon_{g}=\nicefrac{{8}}{{255}}$
@@ 6.3 Ablation and Sensitivity Analysis
We summarize the effect of the hyperparameters $T_{1}$ and $T_{2}$ of Epsilon-Scheduling in Appendix C.1 as follows. (i) $T_{2}$ have the most significant influence with the control of steepness $\left(\nicefrac{{1}}{{(T_{2}-T_{1})}},T_{1}\neq T_{2}\right)$. When $T_{2}$ is close to $T_{1}$, clean accuracy decreases, whereas robust accuracy increases; this eventually leads to suboptimal transfer. This is in line with the motivation for linear warmup in Debenedetti et al. (2023), although they do not study this effect. (ii) Increasing $T_{1}$ increases clean accuracy, up to a threshold beyond which further increasing $T_{1}$ has no apparent effect.
@@
Automated Scheduler Based on our analysis, we can derive a simple automatic epsilon-scheduler (auto) driven by the validation accuracy. The procedure starts with $\epsilon=0$ and then initiates a linear increase from $T_{1}$ to the end of training, where $T_{1}$ is automatically selected as the point at which the validation accuracy converges. Convergence is detected by monitoring the change in validation accuracy with patience of 5 epochs and a tolerance of 2%. Table 4 presents the results obtained with this automatic scheduler, which show that although it has less expected robustness compared to RFT-scheduler, it effectively mitigates suboptimal transfer and provides strong performance across tasks.
$\epsilon=\nicefrac{{4}}{{255}}$ $\epsilon=\nicefrac{{8}}{{255}}$
Model Setting Aircraft Caltech Cars Aircraft Caltech Cars
Clean Adv. E.Adv. Clean Adv. E.Adv. Clean Adv. E.Adv. Clean Adv. E.Adv. Clean Adv. E.Adv. Clean Adv. E.Adv.
auto 73.30 29.40 52.96 85.63 54.18 71.29 84.20 38.40 64.30 68.40 18.60 42.69 81.71 35.82 59.92 79.20 18.70 51.43
CNX fix 7.60 4.50 5.86 83.27 61.54 73.08 69.60 43.20 57.52 1.60 1.50 1.48 59.85 33.95 46.34 5.30 2.60 3.98
sched 78.40 38.00 59.40 89.41 61.45 77.23 88.90 57.70 75.85 75.00 28.80 50.90 84.99 41.82 64.92 85.60 35.90 65.04
auto 79.10 31.60 56.61 90.14 58.30 76.26 89.00 50.30 72.56 76.20 23.60 49.65 86.48 38.39 64.35 86.20 29.90 63.49
Table 4: Results on an automated scheduler derived from our analysis for SWIN and ConvNext (CNX) on Aircraft, Caltech, and Cars, at $\epsilon=\nicefrac{{4}}{{255}}$ (left block) and $\epsilon=\nicefrac{{8}}{{255}}$ (right block).
@@ 7 Conclusion
Limitations and Future Work. Although Epsilon-Scheduling yields significant improvements, robustness can still be limited even when clean accuracy is high, highlighting the potential for future research to further enhance performance. This work opens doors to exploring other scheduling strategies, either heuristic, theoretically motivated, or learning-based. Extending the analysis to other vision tasks, such as detection or segmentation, applying the framework to parameter-efficient methods like LoRA, and investigating whether similar dynamics occur in other modalities, such as natural language processing, remain open questions. Studying these cases may require special considerations such as task-specific losses or hyperparameters.
From a theoretical perspective, although we offer an explanation based on the discrepancy between the clean and robust loss landscapes in the vicinity of the pretrained model, a deeper understanding of robust fine-tuning in this setting remains an open challenge. Our findings point to several important open problems: (i) What mechanisms underlie suboptimal transfer: is delayed task adaptation the only cause of suboptimal transfer or are there other factors? (ii) Can we find other approaches to mitigate delayed task adaptation different from Epsilon-Scheduling? (iii) What mathematical theory can account for suboptimal transfer or delayed task adaptation? (iv) If robust pretraining is not indispensable, what specific properties (if any) in pretraining really matter for downstream robustness and allow effective robust fine-tuning?
Pursuing these directions promises to unlock more effective strategies for robust fine-tuning and yield more substantial progress towards achieving robustness in downstream tasks.